OpenStudy (anonymous):

x=-2+y and -16x-3y=-25

OpenStudy (anonymous):

what are the x and y values

OpenStudy (anonymous):

I am guessing you what to solve for x and y

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hey zoomiemoo what school u go to

OpenStudy (anonymous):

First move the variables to ones side of the equations y-x=2 -16x-3y=-25 Next multiply the first equation by three and add every thing together 3y-3x-16x-3y=6-25

OpenStudy (anonymous):

what are the x and y values though

OpenStudy (anonymous):

Simplify to get -19x=-19 x=1

OpenStudy (anonymous):

next put the x value into one of the equations to find y (1)=-2+y y=3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

finally check if the values works for the second equation -16(1)-3(3)=-25 It works

OpenStudy (anonymous):

it does and thanks

OpenStudy (anonymous):

Your welcome